# Extending Dense Passage Retrieval with Temporal Information

### v1

@@ Abstract.
Temporal awareness is crucial in many information retrieval tasks, particularly in scenarios where the relevance of documents depends on their alignment with the query’s temporal context. Traditional retrieval methods such as BM25 and Dense Passage Retrieval (DPR) excel at capturing lexical and semantic relevance but fall short in addressing time-sensitive queries. To bridge this gap, we introduce the temporal retrieval model that integrates explicit temporal signals by incorporating query timestamps and document dates into the representation space. Our approach ensures that retrieved passages are not only topically relevant but also temporally aligned with user intent. We evaluate our approach on two large-scale benchmark datasets, ArchivalQA and ChroniclingAmericaQA, achieving substantial performance gains over standard retrieval baselines. In particular, our model improves Top-1 retrieval accuracy by 6.63% and NDCG@10 by 3.79% on ArchivalQA, while yielding a 9.56% boost in Top-1 retrieval accuracy and 4.68% in NDCG@10 on ChroniclingAmericaQA. Additionally, we introduce a time-sensitive negative sampling strategy, which refines the model’s ability to distinguish between temporally relevant and irrelevant documents during training. Our findings highlight the importance of explicitly modeling time in retrieval systems and set a new standard for handling temporally grounded queries.
@@ 1. Introduction
In this paper, we explore the following question: Can we develop an effective dense retrieval model that incorporates temporal information alongside the semantic representation of queries and passages? We introduce three novel approaches for integrating temporal aspects into the DPR framework. First, the DateAsTag approach wraps timestamps in special tokens, such as [S-DATE] and [E-DATE], to explicitly mark temporal information in the input. Second, in the DateAsToken approach, we append the timestamp (e.g. year, month, or full date) as plain text to the query and document, treating it as part of the input text. Third, our core contribution, Temporal Dense Passage Retrieval (TempDPR), incorporates temporal embeddings directly into the query and document representations. To achieve this, we explore various fusion techniques, including Feature Stacking (FS), Vector Summation (VS), Relative Embeddings (RE), and Element-Wise Interaction (EWI). These techniques enable the model to encode temporal information within dense embeddings for ensuring robust temporal alignment.
@@
Figure 1. A schematic diagram illustrating an overview of training used in different approaches: (a) VanillaDPR approach, (b) DateAsTag approach, (c) DateAsToken approach, and (d) TempDPR. Due to space limitation we only show here the feature stacking technique for the TempDPR approach. Note that the positive passage’s publication date (PD) may differ from the date in the query, as demonstrated also in the figure’s example.
@@
    Third, we integrate TempDPR into a Retrieval-Augmented Generation (RAG) pipeline. Using large language models (LLMs), we explore the effectiveness of TempDPR’s retrieved results in improving answer generation.
@@ 3. Method
Building on these, we next propose a Temporal Dense Passage Retrieval (TempDPR), which explicitly incorporates both the content and temporal information of queries and passages. Unlike the DateAsTag and DateAsToken approaches, TempDPR represents temporal information as continuous embeddings, enabling it to effectively capture temporal relationships in a learned vector space. To integrate semantic and temporal embeddings, we introduce several fusion techniques, such as feature stacking, vector summation, relative embeddings, and element-wise interaction. These fusion operations allow the model to combine semantic and temporal features in different ways during the retrieval process. All the approaches tested in this paper, including the proposed TempDPR, are illustrated in Figure 1.
@@ 3.1. Temporal Dense Passage Retrieval
where $E_{d}(\cdot)$ encodes the passage content, $E_{t}(\cdot)$ encodes the timestamp, and $\mathrm{Fuse}(\cdot)$ represents the fusion operation used to combine the two embeddings. To jointly model semantic and temporal information, we define three encoding functions: a passage encoder, a query encoder, and a temporal encoder.
The encoding framework in TempDPR consists of components designed to represent both the textual and temporal aspects of queries and passages. A single pretrained BERT model is utilized, producing semantic embeddings for both queries and passages. Specifically, the passage encoder $E_{p}$ maps a passage $p\in\mathcal{P}$ into a semantic embedding space $\mathbb{R}^{d_{s}}$, and the same model is used to map a query $q\in\mathcal{Q}$ into the same embedding space $\mathbb{R}^{d_{s}}$. The final representation is derived from the [CLS] token, where $E_{p}(p)=\text{BERT}(p)_{\texttt{[CLS]}}$ and $E_{q}(q)=\text{BERT}(q)_{\texttt{[CLS]}}$. In addition to semantic embeddings, the temporal encoder $E_{t}$ maps timestamps $t\in\mathcal{T}$ into a temporal embedding space $\mathbb{R}^{d_{t}}$, allowing temporal information to complement the semantic representations. The semantic and temporal embeddings are then integrated using one of four fusion operations to create a unified representation for downstream retrieval tasks.
@@
For TempDPR, we aim to optimize the balance between representing semantic relevance and the ability to capture temporal relevance. Temporal information adds complexity to the retrieval task, as both the content and the time dimension between the query and its matching passages must be aligned. A variety of similarity metrics can be employed for this purpose. We focus on the inner product of the fused semantic and temporal vectors for its simplicity and efficiency. The relevance score between a query $q$ and a passage $p_{i}$ is then computed using the dot product of their encoded vectors:
@@
In TempDPR, the temporal alignment is achieved by embedding timestamps into a shared representation space. Specifically, we encode the query timestamp ($t_{q}$) and passage timestamp ($t_{pi}$) alongside their respective text embeddings, ensuring that temporal signals are weighted equally with semantic content during the similarity computation. We also introduce in Sec. 4.4 temporal negative sampling strategy to address the common issue of temporal misalignment in retrieval tasks, explicitly training the model to distinguish between temporally relevant and irrelevant documents. Consequently, TempDPR can prioritize documents with temporal alignment while maintaining textual relevance. We also note here that the publication date (PD) of a relevant passage does not necessarily need to be the same as the date mentioned in the question. The information relevant to an event is often published around the time of the event, and hence the PD of a positive passage may not always align exactly with the date referenced in the question.
@@ 3.2. Training
In TempDPR model, we employ a BERT-based encoder (Devlin et al., 2019) to capture both the semantic content of the query and passage, as well as their temporal contexts. Unlike conventional retrieval models that rely purely on semantic embeddings, our approach extends the representation by incorporating temporal information using various fusion techniques.
@@
To learn the parameters of the TempDPR retrieval model, we formulate a training objective that explicitly considers both semantic and temporal relevance. The training dataset $\mathcal{D}$, defined as $\mathcal{D}=\{\langle q_{i},t_{q_{i}},p^{+}_{i},t_{p^{+}_{i}},p^{-}_{i,1},t_{p^{-}_{i,1}},\dots,p^{-}_{i,n},t_{p^{-}_{i,n}}\rangle\}_{i=1}^{m}$, where each instance in the dataset contains a query $q_{i}$ and its associated timestamp $t_{q_{i}}$, a positive passage $p^{+}_{i}$ with its timestamp $t_{p^{+}_{i}}$, and $n$ negative passages $\{p^{-}_{i,j},t_{p^{-}_{i,j}}\}$ each with their own timestamps. Our goal is to optimize the retrieval such that the positive passage, which is temporally aligned with the query, is ranked higher than the negative passages. To train the model, we use a cross-entropy loss function that encourages the similarity score between the query and the positive passage to be higher than that between the query and the negative passages while accounting for temporal alignment. Formally, the loss for a single instance is defined as:
@@ 4.1. Temporal Question Answering Datasets
Table 2. Top-{1,5,10,20,50,100} retrieval accuracy for ArchivalQA and ChroniclingAmericaQA datasets. Results are evaluated on both validation and test sets. TempDPR variants consistently outperform state-of-the-art temporal retrieval models.
@@
TempDPR VS 49.23 64.74 70.51 75.14 81.80 85.38 49.72 65.11 71.27 76.72 81.45 84.77
TempDPR RE 48.90 64.34 69.46 75.55 81.48 84.40 47.28 65.67 71.74 75.85 82.24 85.56
TempDPR EWI 47.44 62.88 69.70 74.82 80.18 84.48 46.57 63.38 69.69 75.22 80.66 84.06
TempDPR FS 50.22 64.88 71.16 77.90 83.05 86.22 50.10 66.01 71.85 77.19 82.81 86.52
@@
TempDPR VS 68.42 86.23 90.55 93.30 95.97 97.30 67.96 86.23 90.47 93.55 96.24 97.55
TempDPR RE 66.76 86.93 90.98 94.02 96.40 97.46 66.11 87.13 91.53 94.32 96.83 97.88
TempDPR EWI 69.97 85.40 90.18 93.34 95.71 97.08 69.27 85.79 90.04 93.05 95.85 97.16
TempDPR FS 69.72 87.82 991.38 94.21 96.44 97.77 69.84 87.78 91.65 94.28 96.78 97.75
Table 3. Top-{1,5,10,20,50,100} retrieval accuracy for ArchivalQA and ChroniclingAmericaQA datasets, evaluated on both validation and test sets. The results measure the percentage of top retrieved passages that contain the correct answer across different temporal retrieval models, including the baseline VanillaDPR, DateAsTag, DateAsToken, and TempDPR approaches.
@@
TempDPR FS $\mathbf{69.72}_{\pm 0.34}$ $\mathbf{87.82}_{\pm 0.32}$ $\mathbf{91.38}_{\pm 0.09}$ $\mathbf{94.21}_{\pm 0.06}$ $\mathbf{96.44}_{\pm 0.12}$ $\mathbf{97.77}_{\pm 0.03}$ $\mathbf{69.84}_{\pm 0.86}$ $\mathbf{87.78}_{\pm 0.46}$ $\mathbf{91.65}_{\pm 0.36}$ $\mathbf{94.28}_{\pm 0.35}$ $\mathbf{96.78}_{\pm 0.23}$ $\mathbf{97.75}_{\pm 0.13}$
@@
TempDPR FS $\mathbf{50.22}_{\pm 0.19}$ $\mathbf{64.88}_{\pm 0.23}$ $\mathbf{71.16}_{\pm 0.30}$ $\mathbf{77.90}_{\pm 0.21}$ $\mathbf{83.05}_{\pm 0.11}$ $\mathbf{86.22}_{\pm 0.29}$ $\mathbf{50.10}_{\pm 0.10}$ $\mathbf{66.01}_{\pm 0.17}$ $\mathbf{71.85}_{\pm 0.39}$ $\mathbf{77.19}_{\pm 0.11}$ $\mathbf{82.81}_{\pm 0.25}$ $\mathbf{86.52}_{\pm 0.10}$
@@ 4.3. Selection of Positive Passages
TempDPR FS 44.35 51.79 53.03 54.23 55.24 55.75 44.83 51.44 52.85 53.88 55.01 55.33 44.35 49.78 50.29 50.61 50.77 50.82 44.83 49.52 50.09 50.38 50.57 50.60
Table 5. Performance of BiTimeBERT and TS-Retriever enhanced with TempDPR ’s fusion techniques on ChroniclingAmericaQA and ArchivalQA. Results include Top-1, 5, 10, 20, 50, 100 retrieval accuracy for validation and test sets, demonstrating significant improvements in temporal retrieval performance.
@@
TempDPR +${}_{VS}+\texttt{BiTimeBERT}$ 37.77 57.19 65.15 71.41 78.07 82.45 38.67 57.14 63.14 70.80 77.27 82.79
TempDPR +${}_{RE}+\texttt{BiTimeBERT}$ 35.01 56.13 63.36 69.37 77.66 83.27 37.65 54.70 61.33 68.43 76.32 81.06
TempDPR +${}_{EWI}+\texttt{BiTimeBERT}$ 40.57 56.52 64.69 70.89 77.88 82.45 40.15 56.72 64.14 70.51 77.37 82.56
TempDPR +${}_{FS}+\texttt{BiTimeBERT}$ 40.86 58.57 65.64 72.30 78.47 82.86 41.75 58.96 65.27 70.56 78.69 81.77
@@
TempDPR +${}_{VS}+\texttt{BiTimeBERT}$ 63.70 83.09 88.48 92.12 95.14 96.63 63.88 83.14 88.33 91.96 95.00 96.49
TempDPR +${}_{RE}+\texttt{BiTimeBERT}$ 60.01 80.60 85.84 90.00 93.79 95.89 61.93 80.96 85.93 90.20 93.76 95.72
TempDPR +${}_{EWI}+\texttt{BiTimeBERT}$ 61.28 80.93 86.33 90.19 93.46 95.50 60.52 81.22 86.57 90.76 94.30 95.90
TempDPR +${}_{FS}+\texttt{BiTimeBERT}$ 64.93 84.36 89.04 92.51 95.63 96.95 65.66 84.63 89.64 92.55 95.49 96.81

### v2

# TempRetriever: Fusion-based Temporal Dense Passage Retrieval for Time-Sensitive Questions

@@ Abstract.
Temporal awareness is crucial in many information retrieval tasks, particularly in scenarios where the relevance of documents depends on their alignment with the query’s temporal context. Traditional approaches such as BM25 and Dense Passage Retrieval (DPR) focus on lexical or semantic similarity but tend to neglect the temporal alignment between queries and documents, which is essential for time-sensitive tasks like temporal question answering (TQA). We propose TempRetriever  [^1] The code and the datasets will be available after paper acceptance., a novel extension of DPR that explicitly incorporates temporal information by embedding both the query date and document timestamp into the retrieval process. This allows retrieving passages that are not only contextually relevant but also aligned with the temporal intent of queries. We evaluate TempRetriever on two large-scale datasets—ArchivalQA and ChroniclingAmericaQA—demonstrating its superiority over baseline retrieval models across multiple metrics. TempRetriever achieves a 6.63% improvement in Top-1 retrieval accuracy and a 3.79% improvement in NDCG@10 compared to the standard DPR on ArchivalQA. Similarly, for ChroniclingAmericaQA, TempRetriever exhibits a 9.56% improvement in Top-1 retrieval accuracy and a 4.68% improvement in NDCG@10. We also propose a novel, time-based negative sampling strategy which further enhances retrieval performance by addressing temporal misalignment during training. Our results underline the importance of temporal aspects in dense retrieval systems and establish a new benchmark for time-aware passage retrieval.
@@ 1. Introduction
In this paper, we explore the following question: Can we develop an effective dense retrieval model that incorporates temporal information alongside the semantic representation of queries and passages? We introduce three novel approaches for integrating temporal aspects into the DPR framework. First, the DateAsTag approach wraps timestamps in special tokens, such as [S-DATE] and [E-DATE], to explicitly mark temporal information in the input. Second, in the DateAsToken approach, we append the timestamp (e.g. year, month, or full date) as plain text to the query and document, treating it as part of the input text. Third, our core contribution, Temporal Dense Passage Retrieval (TempRetriever), incorporates temporal embeddings directly into the query and document representations. To achieve this, we explore various fusion techniques, including Feature Stacking (FS), Vector Summation (VS), Relative Embeddings (RE), and Element-Wise Interaction (EWI). These techniques enable the model to encode temporal information within dense embeddings for ensuring robust temporal alignment.
@@
Figure 1. A schematic diagram illustrating an overview of training used in different approaches: (a) VanillaDPR approach, (b) DateAsTag approach, (c) DateAsToken approach, and (d) TempRetriever. Due to space limitation we only show here the feature stacking technique for the TempRetriever approach. Note that the positive passage’s publication date (PD) may differ from the date in the query, as demonstrated also in the figure’s example.
@@
    Third, we integrate TempRetriever into a Retrieval-Augmented Generation (RAG) pipeline. Using large language models (LLMs), we explore the effectiveness of TempRetriever’s retrieved results in improving answer generation.
@@ 3. Method
Building on these, we next propose a Temporal Dense Passage Retrieval (TempRetriever), which explicitly incorporates both the content and temporal information of queries and passages. Unlike the DateAsTag and DateAsToken approaches, TempRetriever represents temporal information as continuous embeddings, enabling it to effectively capture temporal relationships in a learned vector space. To integrate semantic and temporal embeddings, we introduce several fusion techniques, such as feature stacking, vector summation, relative embeddings, and element-wise interaction. These fusion operations allow the model to combine semantic and temporal features in different ways during the retrieval process. All the approaches tested in this paper, including the proposed TempRetriever, are illustrated in Figure 1.
@@ 3.1. Temporal Dense Passage Retrieval
where $E_{p}(\cdot)$ encodes the passage content, $E_{t}(\cdot)$ encodes the timestamp, and $\mathrm{Fuse}(\cdot)$ represents the fusion operation used to combine the two embeddings. To jointly model semantic and temporal information, we define three encoding functions: a passage encoder, a query encoder, and a temporal encoder.
The encoding framework in TempRetriever consists of components designed to represent both the textual and temporal aspects of queries and passages. A single pretrained BERT model is utilized, producing semantic embeddings for both queries and passages. Specifically, the passage encoder $E_{p}$ maps a passage $p\in\mathcal{P}$ into a semantic embedding space $\mathbb{R}^{d_{s}}$, and the same model is used to map a query $q\in\mathcal{Q}$ into the same embedding space $\mathbb{R}^{d_{s}}$. The final representation is derived from the [CLS] token, where $E_{p}(p)=\text{BERT}(p)_{\texttt{[CLS]}}$ and $E_{q}(q)=\text{BERT}(q)_{\texttt{[CLS]}}$. In addition to semantic embeddings, the temporal encoder $E_{t}$ maps timestamps $t\in\mathcal{T}$ into a temporal embedding space $\mathbb{R}^{d_{t}}$, allowing temporal information to complement the semantic representations. The semantic and temporal embeddings are then integrated using one of four fusion operations to create a unified representation for downstream retrieval tasks.
@@
For TempRetriever, we aim to optimize the balance between representing semantic relevance and the ability to capture temporal relevance. Temporal information adds complexity to the retrieval task, as both the content and the time dimension between the query and its matching passages must be aligned. A variety of similarity metrics can be employed for this purpose. We focus on the inner product of the fused semantic and temporal vectors for its simplicity and efficiency. The relevance score between a query $q$ and a passage $p_{i}$ is then computed using the dot product of their encoded vectors:
@@
In TempRetriever, the temporal alignment is achieved by embedding timestamps into a shared representation space. Specifically, we encode the query timestamp ($t_{q}$) and passage timestamp ($t_{pi}$) alongside their respective text embeddings, ensuring that temporal signals are weighted equally with semantic content during the similarity computation. We also introduce in Sec. 4.4 temporal negative sampling strategy to address the common issue of temporal misalignment in retrieval tasks, explicitly training the model to distinguish between temporally relevant and irrelevant documents. Consequently, TempRetriever can prioritize documents with temporal alignment while maintaining textual relevance. We also note here that the publication date (PD) of a relevant passage does not necessarily need to be the same as the date mentioned in the question. The information relevant to an event is often published around the time of the event, and hence the PD of a positive passage may not always align exactly with the date referenced in the question.
@@ 3.2. Training
In TempRetriever model, we employ a BERT-based encoder (Devlin et al., 2019) to capture both the semantic content of the query and passage, as well as their temporal contexts. Unlike conventional retrieval models that rely purely on semantic embeddings, our approach extends the representation by incorporating temporal information using various fusion techniques.
@@
To learn the parameters of the TempRetriever retrieval model, we formulate a training objective that explicitly considers both semantic and temporal relevance. The training dataset $\mathcal{D}$, defined as $\mathcal{D}=\{\langle q_{i},t_{q_{i}},p^{+}_{i},t_{p^{+}_{i}},p^{-}_{i,1},t_{p^{-}_{i,1}},\dots,p^{-}_{i,n},t_{p^{-}_{i,n}}\rangle\}_{i=1}^{m}$, where each instance in the dataset contains a query $q_{i}$ and its associated timestamp $t_{q_{i}}$, a positive passage $p^{+}_{i}$ with its timestamp $t_{p^{+}_{i}}$, and $n$ negative passages $\{p^{-}_{i,j},t_{p^{-}_{i,j}}\}$ each with their own timestamps. Our goal is to optimize the retrieval such that the positive passage, which is temporally aligned with the query, is ranked higher than the negative passages. To train the model, we use a cross-entropy loss function that encourages the similarity score between the query and the positive passage to be higher than that between the query and the negative passages while accounting for temporal alignment. Formally, the loss for a single instance is defined as:
@@ 4.1. Temporal Question Answering Datasets
Table 2. Top-{1,5,10,20,50,100} retrieval accuracy for ArchivalQA and ChroniclingAmericaQA datasets. Results are evaluated on both validation and test sets. TempRetriever variants consistently outperform state-of-the-art temporal retrieval models.
@@
TempRetriever VS 49.23 64.74 70.51 75.14 81.80 85.38 49.72 65.11 71.27 76.72 81.45 84.77
TempRetriever RE 48.90 64.34 69.46 75.55 81.48 84.40 47.28 65.67 71.74 75.85 82.24 85.56
TempRetriever EWI 47.44 62.88 69.70 74.82 80.18 84.48 46.57 63.38 69.69 75.22 80.66 84.06
TempRetriever FS 50.22 64.88 71.16 77.90 83.05 86.22 50.10 66.01 71.85 77.19 82.81 86.52
@@
TempRetriever VS 68.42 86.23 90.55 93.30 95.97 97.30 67.96 86.23 90.47 93.55 96.24 97.55
TempRetriever RE 66.76 86.93 90.98 94.02 96.40 97.46 66.11 87.13 91.53 94.32 96.83 97.88
TempRetriever EWI 69.97 85.40 90.18 93.34 95.71 97.08 69.27 85.79 90.04 93.05 95.85 97.16
TempRetriever FS 69.72 87.82 91.38 94.21 96.44 97.77 69.84 87.78 91.65 94.28 96.78 97.75
Table 3. Top-{1,5,10,20,50,100} retrieval accuracy for ArchivalQA and ChroniclingAmericaQA datasets, evaluated on both validation and test sets. The results measure the percentage of top retrieved passages that contain the correct answer across different temporal retrieval models, including the baseline VanillaDPR, DateAsTag, DateAsToken, and TempRetriever approaches.
@@
TempRetriever FS $\mathbf{69.72}_{\pm 0.34}$ $\mathbf{87.82}_{\pm 0.32}$ $\mathbf{91.38}_{\pm 0.09}$ $\mathbf{94.21}_{\pm 0.06}$ $\mathbf{96.44}_{\pm 0.12}$ $\mathbf{97.77}_{\pm 0.03}$ $\mathbf{69.84}_{\pm 0.86}$ $\mathbf{87.78}_{\pm 0.46}$ $\mathbf{91.65}_{\pm 0.36}$ $\mathbf{94.28}_{\pm 0.35}$ $\mathbf{96.78}_{\pm 0.23}$ $\mathbf{97.75}_{\pm 0.13}$
@@
TempRetriever FS $\mathbf{50.22}_{\pm 0.19}$ $\mathbf{64.88}_{\pm 0.23}$ $\mathbf{71.16}_{\pm 0.30}$ $\mathbf{77.90}_{\pm 0.21}$ $\mathbf{83.05}_{\pm 0.11}$ $\mathbf{86.22}_{\pm 0.29}$ $\mathbf{50.10}_{\pm 0.10}$ $\mathbf{66.01}_{\pm 0.17}$ $\mathbf{71.85}_{\pm 0.39}$ $\mathbf{77.19}_{\pm 0.11}$ $\mathbf{82.81}_{\pm 0.25}$ $\mathbf{86.52}_{\pm 0.10}$
@@ 4.3. Selection of Positive Passages
TempRetriever FS 44.35 51.79 53.03 54.23 55.24 55.75 44.83 51.44 52.85 53.88 55.01 55.33 44.35 49.78 50.29 50.61 50.77 50.82 44.83 49.52 50.09 50.38 50.57 50.60
Table 5. Performance of BiTimeBERT and TS-Retriever enhanced with TempRetriever ’s fusion techniques on ChroniclingAmericaQA and ArchivalQA. Results include Top-1, 5, 10, 20, 50, 100 retrieval accuracy for validation and test sets, demonstrating significant improvements in temporal retrieval performance.
@@
TempRetriever +${}_{VS}+\texttt{BiTimeBERT}$ 37.77 57.19 65.15 71.41 78.07 82.45 38.67 57.14 63.14 70.80 77.27 82.79
TempRetriever +${}_{RE}+\texttt{BiTimeBERT}$ 35.01 56.13 63.36 69.37 77.66 83.27 37.65 54.70 61.33 68.43 76.32 81.06
TempRetriever +${}_{EWI}+\texttt{BiTimeBERT}$ 40.57 56.52 64.69 70.89 77.88 82.45 40.15 56.72 64.14 70.51 77.37 82.56
TempRetriever +${}_{FS}+\texttt{BiTimeBERT}$ 40.86 58.57 65.64 72.30 78.47 82.86 41.75 58.96 65.27 70.56 78.69 81.77
@@
TempRetriever +${}_{VS}+\texttt{BiTimeBERT}$ 63.70 83.09 88.48 92.12 95.14 96.63 63.88 83.14 88.33 91.96 95.00 96.49
TempRetriever +${}_{RE}+\texttt{BiTimeBERT}$ 60.01 80.60 85.84 90.00 93.79 95.89 61.93 80.96 85.93 90.20 93.76 95.72
TempRetriever +${}_{EWI}+\texttt{BiTimeBERT}$ 61.28 80.93 86.33 90.19 93.46 95.50 60.52 81.22 86.57 90.76 94.30 95.90
TempRetriever +${}_{FS}+\texttt{BiTimeBERT}$ 64.93 84.36 89.04 92.51 95.63 96.95 65.66 84.63 89.64 92.55 95.49 96.81